OpenStudy (maheshmeghwal9):

Can anyone teach me how to write a complex no. in its trigonometrical or polar form?

terenzreignz (terenzreignz):

Sure, so what do you know about it so far? :)

OpenStudy (maheshmeghwal9):

I know this much: that if there is a complex number (x+iy) Then its polar form is \[re^{i \theta}.\]where \[r=\sqrt{x^2+y^2}.\]& \[\theta = argument\]& trigonometrical form is \[x+iy=r(\cos \theta+i \sin \theta).\]

OpenStudy (maheshmeghwal9):

I mainly have problem in writing argument of a complex number in its trigonometrical form.

OpenStudy (anonymous):

Writing a complex number of the form x + iy in polar coordinates of the form rcos θ + irsin θ is called Polar Form of the Complex Number.

OpenStudy (anonymous):

A complex number can be written in any of the following polar forms: rcos θ + irsin θ, r(cos θ + isin θ), or rcis θ, where r is known as modulus and θ is known as argument.

OpenStudy (maheshmeghwal9):

i know that but my problem is other:/

OpenStudy (anonymous):

i give you an example Express the complex number - 5 + 5i in polar form.

terenzreignz (terenzreignz):

If your complex number is of the form a + ib (I'll use a and b instead of x and y, because I reserve those for rectangular coordinates, I hope you don't mind) Then your r is as you have written, so there's no problem with that :) As for your argument Take the absolute values of a and b and plug them in here : \[\tan ^{-1}\frac{|b|}{|a|}\] Note the result, we'll just call it your angle :) (can't think of a better word) Now If BOTH a and b were positive, then your argument is precisely your angle :) If a is negative and b is positive, then your argument is 180 minus your angle :) If BOTH a and b were negative, then your argument is 180 plus your angle :) If a is positive and b is negative, then your argument is 360 minus your angle :) so, does that help? :)

OpenStudy (maheshmeghwal9):

yes thank u. My main problem in ur last paragraph is solved but will u do some examples with me?

terenzreignz (terenzreignz):

@nitz gave you one Do you want to do it? :)

OpenStudy (anonymous):

 think of complex number like a point on real plane. And use all trigonometry you know same way like in R2